A football was moving at a speed of 1.2 ms-1 when it was kicked towards the opposite direction.  The impact time was 0.2 s and the final speed was 5.0 m s-1.  Find the average acceleration of the ball during the impact.


1
Expert's answer
2022-05-18T08:32:09-0400

The average acceleration

"a=\frac{v_f-v_i}{\Delta t}=\frac{5.0-(-1.2)}{0.2}=31\:\rm m/s^2"